what would be the formula for the question below
In what ratio must 35 be increased to become 49.

Answer by Theo(13342)

the ratio of 49 to 35 is equal to 49/35 = 7/5

you would multiply 35 by 7/5 to make it become 49.

you would increase 35 by (7/5 - 5/5) = 2/5 to make it become 49.

35 * 7/5 = 49

35 + 35 * 2/5 = 49

7/5 is equal to 1.4

2/5 is equal to .4

35 * 7/5 is the same as 35 * 1.4 which is the same as 35 * (1 + .4) which is equal to 35 + .4 * 35 = 49

35 + 2/5 * 35 is the same as 35 + .4 * 35 which is equal to 35 * (1 + .4) which is equal to 35 * 1.4 = 49

so, ........

if you want to know what ratio you need to multiply 35 by in order to get 49, you would get that ratio as 49/35 = 7/5 = 1.4

if you want to know what ratio you have to add to 35 in order to get 49, you would get that ratio as (49 - 35) / 35 = 14/35 = 2/5 = .4

i believe your question is in what ratio does 35 have to be INCREASED.

that would indicate 2/5 would be the answer as best i can determine.

35 + 2/5 * 35 = 49

when you are dealing with percents, you are faced with the same type of problem.

percent and ratios are related in that percent is equal to 100 * ratio and ratio = percent / 100.

the ratio is also called the decimal equivalent of the percent.

example:

what percent is 100 of 50?

it is 100/50 = 2 (that's the ratio) * 100 = 200% (that's the percent).

what percent is 100 more than 50, or:
what percents does 50 have to be increased by to equal 100?

it is (100 - 50) / 50 = 1 (that's the ratio) * 100 = 100% (that's the percent.

100 is 100% more than 50.
100 is an increase of 100% over 50.
you would have to increase 50 by 100% in order to get 100.